Exhibit 22-2
On January 1, 2015, Nathan, Inc. purchased a machine for $56,000. Eight-year, straight-line depreciation with no salvage value was used through December 31, 2016. On January 1, 2017, it was estimated that the total useful life of the machine from acquisition date was ten years.

Refer to Exhibit 22-2. The adjusting entry that should be made on January 1, 2017, will be in the amount of

$6,000

$3,600

$2,400

$       0
Answer #1

Depreciation for 2 years = $56,000 / 8 * 2 = $14,000

Book value of machine as on January 1, 2017 = $56,000 - 14,000 = $42,000

Revised depreciation expense = $42,000 / 8 = $5,250.

Hence adjusting entry made on January 1, 2017 will be in amount of $0.

Hence option $0 is correct.